Assume a simple model without any government or net exports. If a decrease in autonomous investment by 40 leads to a final decrease in GDP by 160, then the marginal propensity to save is _________. Show your calculations.
Answer #1

Answer

Multiplier =change in GDP/change in investment

=-160/(-40)

=4

Multiplier =1/MPS

MPS=1/multiplier =1/4=0.25

the MPS(marginal propensity to consume) is 0.25